# Asymptotic Velocity Domination

### v1

@@ 2.3 Fourier decomposition and canonical quantization
where the time variable is taken to be $\tau=\ln t/t_{0}$, with $t\in\mathbb{R}_{+}$ transcribing to $\tau\in\mathbb{R}$, and the Big Bang located at $\tau=$-$\infty$. Classically, the $a_{n},a_{n}^{*}$ are just complex parameters; later, they will be promoted to Fock space operators. Applying a spatial Fourier transform to the wave equation for $\phi$ in Eq. (LABEL:evolpt) for $\rho=t$ yields
@@ 2.4 Two-point functions of gauge-fixed Dirac observables
${\cal O}(\theta)$ $\!\!=\!\!$ $\int_{0}^{2\pi}\frac{d\zeta}{2\pi}q(\tau,\zeta;\theta)\,,\quad{\rm Im}\theta\neq 0\,$,
$q(\tau,\zeta;\theta)$ $:=$ $c_{0}(\zeta\!+\!\theta,e^{\tau})\partial_{\tau}\phi+e^{2\tau}\partial_{\zeta}c_{1}(\zeta\!+\!\theta,e^{\tau})\phi\,$, (2.31)
@@ 3.  Time evolution of two-point functions
The time evolution of the Heisenberg picture field operators and Fock space operators can be described by convolution with a symplectic matrix kernel that is defined in terms of the commutator function; see (LABEL:opevol). This in turn gives rise to an initial value parameterization for the symmetric part of the two-point functions; see (LABEL:twopmat2), (LABEL:twopmat3). For the present purposes, the issue whether or not the time evolution is implemented by a unitary operator on Fock space [12, 13] is secondary. For simplicity, we set $\lambda_{{\textsc{n}}}=1$ throughout this section.
@@ 3.1 Time evolution of field operators
The solution of the initial value problem of Heisenberg picture field operators can concisely be expressed in terms of the commutator function $\Delta(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ defined in (2.4). Its Fourier components $\Delta_{n}(\tau,\tau^{\prime})$ are real valued and can be characterized by the following relations: $\partial_{\tau}^{2}+n^{2}e^{2\tau}]\Delta_{n}(\tau,\tau^{\prime})=0$, $\Delta_{n}(\tau,\tau^{\prime})=-\Delta_{n}(\tau^{\prime},\tau)$, $\partial_{\tau}\Delta_{n}(\tau,\tau^{\prime})|_{\tau^{\prime}=\tau}=1$. Merely on account of these properties one has
@@ 3.2 Initial value parameterization and time consistent states
Applied to (LABEL:twopmat2) this means that two initial value kernels ${\cal W}^{s}(\tau_{0},\tau_{0},\,\cdot\,)$ and ${\cal W}^{s}(\tau_{1},\tau_{1},\,\cdot\,)$ must be related by the $\tau=\tau^{\prime}=\tau_{1}$ specialization of (LABEL:twopmat2). In the Fourier representation (LABEL:twopmat3) this amounts to
@@
Result. The transformations (2.25), (2.30) trivialize the consistency conditions (3.17), (3.18) for the initial data under a change of reference time: the latter hold iff the Bogoliubov parameters $\lambda_{n}(\tau_{0}),\mu_{n}(\tau_{0})$ and $\tilde{\lambda}_{n}(\tau_{0}),\tilde{\mu}_{n}(\tau_{0})$, $n\in\mathbb{Z}$, are independent of $\tau_{0}$. Further, both sets of Bogoliubov parameters are consistently related by the time independent relation (3.21) below. Both sets of time consistent initial data are in turn related by $I^{\rm grad}(|n|e^{\tau_{0}})$ from (3.2). Overall, the four maps give rise to the commutative diagram in Figure 2.
@@ 3.3 Spatial gradient expansion of two-point functions
In position space, the velocity dominated two-point functions will in general contain distributional terms in $\zeta-\zeta^{\prime}$ and will not be conventionally differentiable to all orders. However once spatially averaged with real, smooth test functions $f,g\in C^{\infty}(S^{1})$, the differentiations can be defined distributionally. That is, for any element $\mathfrak{F}$ in the dual [^5] That is, $\mathfrak{F}(\zeta)=\sum_{m}\mathfrak{F}_{m}e^{im\zeta}$, where there exists $c^{\prime}>0,M\in\mathbb{N}$ such that $|\mathfrak{F}_{m}|\leq c^{\prime}(1+|m|)^{M}$, for all $m\in\mathbb{Z}$. This is is dual to the characterization of $f\in C^{\infty}(S^{1})$, i.e. $f(\zeta)=\sum_{n}f_{n}e^{in\zeta}$, with $|f_{n}|\leq c(1+|n|)^{N}$, for all $N\in\mathbb{N}$. of $C^{\infty}(S^{1})$ and $l\in\mathbb{N}$ we set
@@
Result. The matrix two-point function ${\cal W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ admits upon averaging with $f,g\in C^{\infty}(S^{1})$ test functions a series expansion of the form
@@
Here, $I_{k}$ are the numerical $2\times 2$ matrices in the expansion of $I^{\rm grad}(x)$ in powers of $x^{2}$. All terms in the expansion are spatial gradients of the two-point function $\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ of the velocity dominated system, which is linear in $\tau,\tau^{\prime}$. The series (1) is uniformly convergent for all $\tau,\tau^{\prime}<-\delta$, $\delta>0$ and $\tau-\tau^{\prime}$ bounded.
Only the convergence part needs to be shown. This can be done using the Peano form of the Taylor series remainder (applied to $I^{\rm grad}(x)$) in combination with matrix norm bounds of the form entering in Section 3.4. We omit the details.
@@ 3.4 Asymptotic velocity domination for two-point functions
The $n=0$ term drops out on account of zero modes identification. Next, we insert (3.32) and split off the unit matrix part from $I^{\rm{grad}}$, defining $\hat{I}(x)=I^{\rm{grad}}(x)-\mbox{1\hskip-3.44444pt1}$. Termwise this gives
@@
$\quad+\|\hat{I}(|n|e^{\tau})\|_{\infty}\|\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})\|_{\infty}\|\hat{I}(|n|e^{\tau^{\prime}})^{T}\|_{\infty}\Big{\}}.$ (3.41)
@@
This is still $\tau_{0}$ dependent, but for time consistent initial data we know that $\tilde{z}_{n}(\tau_{0}),\tilde{w}_{n}(\tau_{0})$ arise from $\tau_{0}$-independent Bogoliubov parameters $\tilde{\lambda}_{n},\tilde{\mu}_{n}$ via (2.30). Inserted into (3.45) one has
@@
Result. The difference between the spatially averaged matrix two-point function ${\cal W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$ in the full Gowdy system and its counterpart $\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$ in the velocity dominated system can for $f,g\in C^{\infty}(S^{1})$ be bounded as follows:
@@
(i) This bound entails the desired result: if $\tau,\tau^{\prime},\tau_{0}$ are back-propagated towards the Big Bang via $\tau=\tilde{\tau}+\eta$, $\tau^{\prime}=\tilde{\tau}^{\prime}+\eta$, $\tau_{0}=\tilde{\tau}_{0}+\eta$, $\eta\rightarrow$-$\infty$, the right hand side vanishes like $O(\eta^{2}e^{2\eta})$. This also explains the rationale for the domain of uniform convergence for the series (LABEL:grad20).
@@
For illustration we show in Fig. 3 the actual behavior of the difference for the two-point functions in the Bunch-Davies vacuum, as detailed in section 4.
@@ 4.1 Hadamard condition in Fourier space
Result[25]. Let $T_{p}$ be a Wronskian-normalized solution of $\bigl{[}\partial_{\tau}^{2}+\nu(\tau)+|p|^{2}w(\tau)\bigr{]}T_{p}(\tau)=0$, with $\nu,w\in C^{\infty}(\mathbb{R})$, $w>0$, and $p=(p_{1},\ldots,p_{d})$ the spatial momentum. Define the two-point function
@@
$G_{1}$ $\!=\!$ $\frac{v}{2w}+\frac{5}{32}\frac{{w^{\prime}}^{2}}{w^{3}}-\frac{1}{8}\frac{w^{\prime\prime}}{w^{2}}\,$,
$G_{2}$ $\!=\!$ $\frac{3}{8w^{2}}\Big{(}v^{2}+\frac{1}{3}v^{\prime\prime}\Big{)}-\frac{5}{16w^{3}}\Big{(}vw^{\prime\prime}+v^{\prime}w^{\prime}-v\frac{7{w^{\prime}}^{2}}{4w}\Big{)}$ (4.3)
@@
The asymptotics (4.2) is consistent with the one induced by the adiabatic iteration but much streamlined, see [26] and Appendix A of [25]. Heuristically, Hadamard states are adiabatic vacua of infinite order. Mathematically, this is expressed by Lemma III.2 of [25] which we assume to extend to generic $d\geq 1$. Importantly, the Fourier space Hadamard condition (4.2) only prescribes the large $|p|$ asymptotics of $|T_{p}(\tau)|^{2}$. Different Hadamard states will differ by their non-universal small $|p|$ behavior. It is commonly assumed that a state is also infrared finite, which requires $|T_{p}(\tau)|^{2}=o(|p|^{-\nu})$, $\nu<d$, for the integral to be infrared finite. For example, the States of Low Energy are Hadamard states with $\nu=1$ for all $d\geq 1$, so they are infrared finite for $d\geq 2$ [25].
@@
where the coefficients turn out to be those of $(\pi/4)|H_{0}^{(2)}(|n|e^{\tau})|^{2}$. Here we transitioned to discrete momenta and circular spatial sections. This is expected to be legitimate on general grounds, but in the situation at hand follows more directly from an identity of the form
@@
Here $W^{s}_{\mathbb{R}}$ is the symmetric two-point function for noncompact spatial sections and the term in square brackets is the previous two-point function $W^{s}$ from (LABEL:Wsdef) defined on $S^{1}$, with the zero mode term $W_{0}(\tau,\tau^{\prime})$ removed. The function $c_{N}(\tau,\tau^{\prime})$ is explicitly computable and is regular at $\tau=\tau^{\prime}$. It diverges for $N\rightarrow\infty$ in a way that depends on the infrared properties of $W^{s}_{\mathbb{R}}$’s Fourier transform. In $d=1$ one needs this Fourier transform to scale as $p^{-\nu},0\leq\nu<1$, for the integral to be infrared finite. In this case the leading divergence is $O(N^{\nu}\ln^{2}N)$. For $\nu=1$, relevant for States of Low Energy, the Fourier integral needs to be cut off at some small $0<\mu/(2N+1)\ll 1$. In this case the leading divergence is $O(N\ln^{2}\mu)$ or $O(N\ln^{2}N\ln\mu)$. Note that the periodic two-point function is always infrared finite as $W_{0}(\tau,\tau^{\prime})$ of the form in (LABEL:Wsdef) can be chosen at will. Based on (4.5) one can recover $W^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})-W_{0}(\tau,\tau^{\prime})$ as a subtracted limit of $2\pi W_{\mathbb{R}}^{s}$’s periodic extension. The notion of a Hadamard state thereby carries over from the noncompact to the compact spatial sections without having to construct the parametrix in the latter case. We now discuss two important instances of so-understood Hadamard states for the $T^{3}$-Gowdy system.
@@
This can be recognized as the wave equation of a free massive field on $1+1$ dimensional de Sitter space in the Poincaré patch and conformal time $t$, see e.g. [27]. The mass parameter here is $m=1/2$, which is at the unitarity threshold. For noncompact spatial sections the two-point function $W_{\mathbb{R}}(t,t^{\prime},\zeta\!-\!\zeta^{\prime})$ can be computed explicitly and indeed coincides with the known $W_{{\rm dS}_{2}}(t,t^{\prime},\zeta\!-\!\zeta^{\prime})/\sqrt{tt^{\prime}}$, expressible in terms of an $_{2}F_{1}$ hypergeometric function for $m=1/2$ and with an appropriate $i\epsilon$-prescription. The symmetric part can be expressed as follows
@@
Here ${\bf K}(z)=\int_{0}^{\pi/2}\!ds(1-z\sin s)^{-1/2}$ is the complete elliptic integral and $\xi=[(\zeta-\zeta^{\prime})^{2}-(t-t^{\prime})^{2}]/(4tt^{\prime})$ is the de Sitter embedding distance. The relation (4.5) can be directly verified in this case and used to (re-)construct the periodic two-point function (LABEL:Wsdef) with the above Bogoliubov parameters.
@@
Here ${\cal E}_{n}^{\rm SLE}$ is the energy of the time averaged Hamiltonian in the SLE. By inverting the $c_{1,n},c_{2,n},c_{3,n}$ versus $\lambda_{n},\mu_{n}$ relations in (LABEL:GDGowdysol) one obtains expressions for the $f$-dependent Bogoliubov parameters $\lambda_{n}^{f},\mu_{n}^{f}$ that define the SLE solution $T_{n}^{\rm SLE}(\tau)$ in (2.3). The induced $n$-dependence is rather complicated and it is not obvious that $|T_{n}^{\rm SLE}(\tau)|^{2}$ will satisfy (4.4). Actually, its does as one can show
@@
In summary, both the Bunch-Davies state and the SLE define Hadamard states for the Gowdy scalar by specifying a set of Bogoliubov parameters in (2.3). Via (2.25) the Bogoliubov parameters define initial data, which enter the initial value parameterization of the two-point function (LABEL:twopmat2), (LABEL:twopmat3). The result in Section 3.2 underlying Figure 2 delineates conditions under which these states are also time consistent.

### v2

@@ 2.3 Fourier decomposition and canonical quantization
where the time variable is taken to be $\tau=\ln t/t_{0}$, with $t\in\mathbb{R}_{+}$ transcribing to $\tau\in\mathbb{R}$, and the Big Bang located at $\tau=$-$\infty$. Classically, the $a_{n},a_{n}^{*}$ are just complex parameters; later, they will be promoted to Fock space operators, see [11, 12, 13, 14]. Applying a spatial Fourier transform to the wave equation for $\phi$ in Eq. (LABEL:evolpt) for $\rho=t$ yields
@@ 2.4 Two-point functions of gauge-fixed Dirac observables
${\cal O}(\theta)\!$ $\!\!=\!\!$ $\!\int_{0}^{2\pi}\frac{d\zeta}{2\pi}q(\tau,\zeta;\theta)\,,\quad{\rm Im}\theta\neq 0\,$, (2.32)
$q(\tau,\zeta;\theta)\!$ $\!\!:=\!\!$ $\!c_{0}(\zeta\!+\!\theta,e^{\tau})\partial_{\tau}\phi+e^{2\tau}\partial_{\zeta}c_{1}(\zeta\!+\!\theta,e^{\tau})\phi=\jmath_{0}^{{\textsc{p}}}(e^{\tau},\zeta;\theta)+\partial_{\zeta}[e^{2\tau}c_{1}(\zeta\!+\!\theta,e^{\tau})\phi]\,$,
@@ 3.  Time evolution of two-point functions
The time evolution of the Heisenberg picture field operators and Fock space operators can be described by convolution with a symplectic matrix kernel that is defined in terms of the commutator function; see (LABEL:opevol). This in turn gives rise to an initial value parameterization for the symmetric part of the two-point functions; see (LABEL:twopmat2), (LABEL:twopmat3). The bonus property that the time evolution of (slightly modified) Heisenberg picture field operators can be unitarily implemented on Fock space [13] will not be needed. For simplicity, we set $\lambda_{{\textsc{n}}}=1$ throughout this section.
@@ 3.1 Time evolution of field operators
The solution of the initial value problem of Heisenberg picture field operators can concisely be expressed in terms of the commutator function $\Delta(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ defined in (2.4). Its Fourier components $\Delta_{n}(\tau,\tau^{\prime})$ are real valued and can be characterized by the following relations: $\partial_{\tau}^{2}+n^{2}e^{2\tau}]\Delta_{n}(\tau,\tau^{\prime})=0$, $\Delta_{n}(\tau,\tau^{\prime})=-\Delta_{n}(\tau^{\prime},\tau)$, $\partial_{\tau}\Delta_{n}(\tau,\tau^{\prime})|_{\tau^{\prime}=\tau}=~1$. Merely on account of these properties one has
@@ 3.2 Initial value parameterization and time consistent states
The definition draws on the fact that a homogeneous pure state in a linear quantum field theoretical system can be characterized by the spatial Fourier transform of an inital data matrix [24]. The requirement that the two-point function associated with such a state is independent of the time $\tau_{0}$ where the initial data are specified is a very natural requirement, as the standard definition of a two-point function in terms of the expectation value of a product of smeared field operators does not refer to $\tau_{0}$. To the best of our knowledge it has never been formulated as an explicit requirement, since in earlier uses of the initial value formulation $\tau_{0}$ was always kept fixed [24, 31, 21]. In the present context often the initial data of intermediary or derived quantities ought to be backpropagated to the Big Bang as well, so $\tau_{0}$ cannot be kept fixed. In the following we characterize the initial data that give rise to a reconstructed two-point function independent of $\tau_{0}$. For definiteness we do so directly for the Gowdy scalar at hand, i.e. (LABEL:twopmat3) versus (LABEL:twopmat2), (2.4), but the ensued Result 3.1 would generalize to other linear quantum field theories on spatially homogeneous backgrounds. In Section 4.2 we shall see that these initial data are typically compatible with the Hadamard property.
We begin by spelling out the consistency condition imposed on $Z_{n}(\tau_{0})$ by time consistency. In view of (LABEL:twopmat2) this means that two initial value kernels ${\cal W}^{s}(\tau_{0},\tau_{0},\,\cdot\,)$ and ${\cal W}^{s}(\tau_{1},\tau_{1},\,\cdot\,)$ must be related by the $\tau=\tau^{\prime}=\tau_{1}$ specialization of (LABEL:twopmat2). In the Fourier representation (LABEL:twopmat3) this amounts to
@@
Result 3.1.
The transformations (2.26), (2.31) trivialize the consistency conditions (3.17), (3.18) for the initial data under a change of reference time: the latter hold iff the Bogoliubov parameters $\lambda_{n}(\tau_{0}),\mu_{n}(\tau_{0})$ and $\tilde{\lambda}_{n}(\tau_{0}),\tilde{\mu}_{n}(\tau_{0})$, $n\in\mathbb{Z}$, are independent of $\tau_{0}$. Further, both sets of Bogoliubov parameters are consistently related by the time independent relation (3.21) below. Both sets of time consistent initial data are in turn related by $I^{\rm grad}(|n|e^{\tau_{0}})$ from (3.2). Overall, the four maps give rise to the commutative diagram in Figure 2.
@@ 3.3 Spatial gradient expansion of two-point functions
In position space, the velocity dominated two-point functions will in general contain distributional terms in $\zeta-\zeta^{\prime}$ and will not be conventionally differentiable to all orders. However once spatially averaged with real, smooth test functions $f,g\in C^{\infty}(S^{1})$, the differentiations can be defined distributionally. That is, for any element $\mathfrak{F}$ in the dual [^5] That is, $\mathfrak{F}(\zeta)=\sum_{m}\mathfrak{F}_{m}e^{im\zeta}$, where there exists $c^{\prime}>0,M\in\mathbb{N}$ such that $|\mathfrak{F}_{m}|\leq c^{\prime}(1+|m|)^{M}$, for all $m\in\mathbb{Z}$. This is is dual to the characterization of $f\in C^{\infty}(S^{1})$, i.e. $f(\zeta)=\sum_{n}f_{n}e^{in\zeta}$, with $|f_{n}|\leq c(1+|n|)^{-N}$, for all $N\in\mathbb{N}$. of $C^{\infty}(S^{1})$ and $l\in\mathbb{N}$ we set
@@
Result 3.2.
The matrix two-point function ${\cal W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ admits upon averaging with $f,g\in C^{\infty}(S^{1})$ test functions a series expansion of the form
@@
Here, $I_{k}$ are the numerical $2\times 2$ matrices in the expansion of $I^{\rm grad}(x)$ in powers of $x^{2}$. All terms in the expansion are spatial gradients of the two-point function $\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ of the velocity dominated system, which is linear in $\tau,\tau^{\prime}$. The series (LABEL:grad20) is uniformly convergent for all $\tau,\tau^{\prime}<-\delta$, $\delta>0$ and $|\tau-\tau^{\prime}|$ bounded.
Only the convergence part needs to be shown. We start from (3.32) and insert the Taylor series $I^{\rm grad}(x)=\sum_{k=0}^{K}I_{k}x^{2k}+C_{K}x^{2K}$ with Peano remainder, i.e. $C_{K}$ goes to zero as $K\rightarrow\infty$ in one (and hence any) $2\times 2$ matrix norm. A brief computation gives
${\cal W}^{s}_{n}(\tau,\tau^{\prime})-\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})-\sum_{l=1}^{2K}n^{2l}\Big(\sum_{k=0}^{l}I_{k}\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})I_{l-k}^{T}e^{2k\tau}e^{2(l-k)\tau^{\prime}}\Big)$
$\quad=\sum_{k=0}^{K}n^{2(k+K)}\Big(e^{2k\tau+2K\tau^{\prime}}I_{k}\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})C_{K}^{T}+e^{2k\tau^{\prime}+2K\tau}C_{K}\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})I_{k}^{T}\Big)$
$\quad+n^{4K}e^{2K(\tau+\tau^{\prime})}C_{K}\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})C_{K}^{T}\,.$ (3.39)
In view of (3.3) we multiply this by $f_{n}g_{-n}+f_{-n}g_{n}$ and sum over $n\in\mathbb{N}$. The resulting right hand side is denoted by RHS and we seek to bound it in some $2\times 2$ matrix norm. For definiteness we use the $\|\cdot\|_{\infty}$ norm from (3.40) below. We only sketch the rest of the argument, deferring some of the details to Section 3.4: For large enough $K$ we may assume $\|C_{K}\|_{\infty}\leq 1$. By inspection of (3.2) one sees that $\sum_{k=0}^{K}\|I_{k}\|_{\infty}\leq 2$, for all $K$. This readily gives $\|{\rm RHS}\|_{\infty}\leq 2(2e^{2K\tau}+2e^{2K\tau^{\prime}}+e^{2K(\tau+\tau^{\prime}})\sum_{n\geq 1}|f_{n}||g_{n}|n^{4K}\|\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})\|_{\infty}$. A bound for $\|\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})\|_{\infty}$ will be provided in (3.45), (3.47). Since $|\tilde{\lambda}_{n}|,|\tilde{\mu}_{n}|$ grow like $\log n$ for the relevant cases, see Section 4.2, we may safely bound the sum by $\sum_{n\geq 1}|f_{n}||g_{n}|n^{4K+2\nu}$, for a small power $0<\nu<1$. From footnote 5 we know $|f_{n}|,|g_{n}|\leq c(1+|n|)^{-N}$, for all $N\geq 1$. Taking $N=2\nu+5K$, the previous sum can be bounded by $c\sum_{n\geq 1}n^{-K}=c\,\zeta(K)$. For large $K$ the $\zeta$ function approaches $1$, and for real $K\geq 2$ it is bounded by $2$. It remains to take into account the $\tau,\tau^{\prime}$ powers from (3.45). They lead to terms of the form $e^{2K\tau},e^{2K\tau^{\prime}}$, $\tau e^{2K\tau},\tau^{\prime}e^{2K\tau^{\prime}}$, and $\tau^{\prime}e^{2K\tau},\tau e^{2K\tau^{\prime}}$, $K\geq 2$. All of them are readily seen to be uniformly bounded for $|\tau\!-\!\tau^{\prime}|<d$ and $\tau,\tau^{\prime}<-\delta$, for some (small) $\delta>0$ and (large) $d>0$. This establishes Result 3.2.
@@ 3.4 Asymptotic velocity domination for two-point functions
The $n=0$ term drops out on account of identification of zero modes. Next, we insert (3.32) and split off the unit matrix part from $I^{\rm{grad}}$, defining $\hat{I}(x)=I^{\rm{grad}}(x)-\mbox{1\hskip-3.44444pt1}$. Termwise this gives
@@
$\qquad+\|\hat{I}(|n|e^{\tau})\|_{\infty}\|\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})\|_{\infty}\|\hat{I}(|n|e^{\tau^{\prime}})^{T}\|_{\infty}\Big\}.$ (3.42)
@@
This is still $\tau_{0}$ dependent, but for time consistent initial data we know that $\tilde{z}_{n}(\tau_{0}),\tilde{w}_{n}(\tau_{0})$ arise from $\tau_{0}$-independent Bogoliubov parameters $\tilde{\lambda}_{n},\tilde{\mu}_{n}$ via (2.31). This gives
@@
Result 3.3.
The difference between the spatially averaged matrix two-point function ${\cal W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$ in the full Gowdy system and its counterpart $\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$ in the velocity dominated system can for $f,g\in C^{\infty}(S^{1})$ be bounded as follows:
@@
(i) This bound entails the desired result: if $\tau,\tau^{\prime},\tau_{0}$ are back-propagated towards the Big Bang via $\tau=\tilde{\tau}+\eta$, $\tau^{\prime}=\tilde{\tau}^{\prime}+\eta$, $\tau_{0}=\tilde{\tau}_{0}+\eta$, $\eta\rightarrow$-$\infty$, the right hand side vanishes like $O(\eta^{2}e^{2\eta})$. In particular $\tau-\tau^{\prime}=\tilde{\tau}-\tilde{\tau}^{\prime}$ stays fixed in the limit, mirroring the condition for uniform convergence in Result 3.2. In the latter case the time consistency of the underlying state does not directly enter, as it is tacit on both sides of (3.32).
@@
For illustration, we show in Fig. 3 the actual behavior of the difference for the two-point functions in the Bunch-Davies vacuum, as detailed in Section 4.
@@ 4.1 Hadamard condition in Fourier space
Result[27]. Let $T_{p}$ be a Wronskian‑normalized solution of $\bigl[\partial_{\tau}^{2}+\nu(\tau)+|p|^{2}w(\tau)\bigr]T_{p}(\tau)=0$, with $\nu,w\in C^{\infty}(\mathbb{R})$, $w>0$, and $p=(p_{1},\ldots,p_{d})$ the spatial momentum. Define the two‑point function
@@
$G_{1}$ $\!=\!$ $\frac{\nu}{2w}+\frac{5}{32}\frac{{w^{\prime}}^{2}}{w^{3}}-\frac{1}{8}\frac{w^{\prime\prime}}{w^{2}}\,$,
$G_{2}$ $\!=\!$ $\frac{3}{8w^{2}}\Big(\nu^{2}+\frac{1}{3}\nu^{\prime\prime}\Big)-\frac{5}{16w^{3}}\Big(\nu w^{\prime\prime}+\nu^{\prime}w^{\prime}-\nu\frac{7{w^{\prime}}^{2}}{4w}\Big)$ (4.3)
@@
The asymptotics (4.2) is consistent with the one induced by the adiabatic iteration but much streamlined, see [28, 29] and Appendix A of [27]. Heuristically, Hadamard states are adiabatic vacua of infinite order. Mathematically, this is expressed by Lemma III.2 of [27] which we assume to extend to generic $d\geq 1$. Importantly, the Fourier space Hadamard condition (4.2) only prescribes the large $|p|$ asymptotics of $|T_{p}(\tau)|^{2}$. Different Hadamard states will differ by their non-universal small $|p|$ behavior. It is commonly assumed that a state is also infrared finite, which requires $|T_{p}(\tau)|^{2}=o(|p|^{-\delta})$, $\delta<d$, for the integral to be infrared finite. For example, the States of Low Energy are Hadamard states with $\delta=1$ for all $d\geq 1$, so they are infrared finite for $d\geq 2$ [27].
@@
where the coefficients turn out to be those of $(\pi/4)|H_{0}^{(2)}(|n|e^{\tau})|^{2}$. Here we transitioned to discrete momenta and circular spatial sections. This is expected to be legitimate on general grounds, but in the situation at hand it follows more directly from an identity of the form
@@
Here $W^{s}_{\mathbb{R}}$ is the symmetric two-point function for noncompact spatial sections and the term in square brackets is the previous two-point function $W^{s}$ from (LABEL:Wsdef) defined on $S^{1}$, with the zero mode term $W_{0}(\tau,\tau^{\prime})$ removed. The function $c_{N}(\tau,\tau^{\prime})$ is explicitly computable and is regular at $\tau=\tau^{\prime}$. It diverges for $N\rightarrow\infty$ in a way that depends on the infrared properties of $W^{s}_{\mathbb{R}}$’s Fourier transform. In $d=1$ one needs this Fourier transform to scale as $p^{-\delta},0\leq\delta<1$, for the integral to be infrared finite. In this case the leading divergence is $O(N^{\delta}\ln^{2}N)$. For $\delta=1$, relevant for States of Low Energy, the Fourier integral needs to be cut off at some small $0<\mu/(2N+1)\ll 1$. In this case the leading divergence is $O(N\ln^{2}\mu)$ or $O(N\ln^{2}N\ln\mu)$. Note that the periodic two-point function is always infrared finite as $W_{0}(\tau,\tau^{\prime})$ of the form in (LABEL:Wsdef) can be chosen at will. Based on (4.5) one can recover $W^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})-W_{0}(\tau,\tau^{\prime})$ as a subtracted limit of $2\pi W_{\mathbb{R}}^{s}$’s periodic extension. The notion of a Hadamard state thereby carries over from the noncompact to the compact spatial sections without having to construct the parametrix in the latter case. We now discuss two important instances of so-understood Hadamard states for the $T^{3}$-Gowdy system.
@@
This can be recognized as the wave equation of a free massive field on $1+1$ dimensional de Sitter space in the Poincaré patch and conformal time $t$, see e.g. [30]. Except for the late time limit $t\rightarrow\infty$ it is not conformally invariant. The mass parameter here is $m=1/2$, which is at the unitarity threshold. For noncompact spatial sections the two-point function $W_{\mathbb{R}}(t,t^{\prime},\zeta\!-\!\zeta^{\prime})$ can be computed explicitly and indeed coincides with the known $W_{{\rm dS}_{2}}(t,t^{\prime},\zeta\!-\!\zeta^{\prime})/\sqrt{tt^{\prime}}$, expressible in terms of an $_{2}F_{1}$ hypergeometric function for $m=1/2$ and with an appropriate $i\epsilon$-prescription. The symmetric part can be expressed as follows
@@
Here ${\bf K}(z)=\int_{0}^{\pi/2}\!ds(1-z\sin s)^{-1/2}$ is the complete elliptic integral and $\xi=[(\zeta-\zeta^{\prime})^{2}-(t-t^{\prime})^{2}]/(4tt^{\prime})$ is the de Sitter embedding distance. Near the lightcone $|t-t^{\prime}|=|\zeta-\zeta^{\prime}|$ the function $W^{s}_{\mathbb{R}}$ has a logarithmic singularity, at $t+t^{\prime}=|\zeta-\zeta^{\prime}|$ it is regular, and for $|t-t^{\prime}|\rightarrow\infty$ it behaves like $1/|t-t^{\prime}|$. The relation (4.5) can be directly verified in this case and used to (re-)construct the periodic two-point function (LABEL:Wsdef) with the above Bogoliubov parameters.
@@
Here ${\cal E}_{n}^{\rm SLE}$ is the energy of the time averaged Hamiltonian in the SLE. By inverting the $c_{1,n},c_{2,n},c_{3,n}$ versus $\lambda_{n},\mu_{n}$ relations in (LABEL:GDGowdysol) one obtains expressions for the $f$-dependent Bogoliubov parameters $\lambda_{n}^{f},\mu_{n}^{f}$ that define the SLE solution $T_{n}^{\rm SLE}(\tau)$ in (2.3). The induced $n$-dependence is rather complicated and it is not obvious that $|T_{n}^{\rm SLE}(\tau)|^{2}$ will satisfy (4.4). Actually, it does, as one can show
@@
In summary, both the Bunch-Davies state and the SLE define Hadamard states for the Gowdy scalar by specifying a set of Bogoliubov parameters in (2.3). Via (2.26) the Bogoliubov parameters define initial data, which enter the initial value parameterization of the two-point function (LABEL:twopmat2), (LABEL:twopmat3). The Result 3.1 delineates conditions under which these states are also time consistent.